# LANTERN: Accelerating Visual Autoregressive Models with Relaxed Speculative Decoding

### v1

@@ Appendix B Algorithm
Appendix C Experimental Details
To train text-conditional model’s drafter, we sampled 100k images in LAION-COCO dataset (Chuhmann et al., 2022), which is used to train stage-1 target model. We used same amount of image sampled in ImageNet (Deng et al., 2009) dataset to train class-conditional model’s drafter. Since EAGLE (Li et al., 2024a) reports that benefits of using target model generated sequence as a training data is marginal, we do not use target model generated sequence as a training data to reduce overhead. We used a single-layer decoder with the same structure as the target model, in the same manner of EAGLE. During training, 5% of data is set to be hold out validation dataset.
Since LlamaGen (Sun et al., 2024) use classifier-free guidance (Ho & Salimans, 2021) to generate image, we trained our drafter to both learn conditioned input and null-conditioned input. To do so, we dropped 10% of conditional embedding during trainig, as same as target model training. Batch size is 16 and base learning rate is $10^{-4}$. AdamW (Loshchilov & Hutter, 2019) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ is used and Linear learning rate scheduling with warm up is used with 2000 warm up steps. We select best performing model in terms of top-3 accuracy in hold out validation set for 20 epochs. In addition, Flan-T5 XL (Chung et al., 2022) is used to encode input text for text-conditional generation.
All text-conditional images are generated using a classifier-free guidance scale of 7.5, with top-p set to 0 and top-k set to 1000, which is the default generation configuration of LlamaGen (Sun et al., 2024) official implementation for text-conditional image generation. For class-conditional generation, the classifier-free guidance scale is set to 4.0, with the top-k sampling covering the entire vocabulary and top-p sampling set to 1.0. For EAGLE-2 and our method, 60 candidate tokens are passed into target model for each verification process.
Appendix D Qualitative Results
@@
Figure 7: Qualitative sample for the changes in the generated images according to various $\delta\in\{0.05,0.1,0.2,0.4\}$ and $k\in\{100,300,1000\}$ at $\tau=0$. Input prompt is ’A motorcycle parked in a parking space next to another motorcycle.’. Target model is LlamaGen-XL stage I.
Figure 8: Additional qualitative samples with various $k$ and $\delta$ at $\tau=1$. Target model is LlamaGen-XL stage I and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.
Figure 9: Additional qualitative samples with various $k$ and $\delta$ at $\tau=0$. Target model is LlamaGen-XL stage I and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.
Figure 10: Qualitative sample for the changes in the generated images according to various $\delta\in\{0.05,0.1,0.2,0.4\}$ and $k\in\{100,300,1000\}$ at $\tau=1$. Input prompt is ’A pile of oranges in crates topped with yellow bannanas.’ Target model is LlamaGen-XL stage II.
Figure 11: Additional qualitative samples with various $k$ and $\delta$ at $\tau=1$. Target model is LlamaGen-XL stage II and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.

### v3

@@ D.1 Speculative Decoding with LANTERN
D.2 Proximity Set Construction
Algorithm 2 Proximity Set Construction for LANTERN
1:Latent space representation of tokens, number of neighbors $k$, $D_{TV}$ tolerance $\delta$
2:Precompute $B_{k}(\widetilde{x})$ for all tokens:
3:for each token $\widetilde{x}$ in the quantized latent space do
4:     Compute distances between $\widetilde{x}$ and all other tokens.
5:     Identify $k$ nearest tokens (including $\widetilde{x}$ itself) based on $\ell_{2}$ distance.
6:     Store these tokens as the set $B_{k}(\widetilde{x})$.
7:end for
8:Dynamically calculate $A_{k,\delta}(\widetilde{x})$ during inference:
9:for each token $\widetilde{x}$ sampled during decoding do
10:     Initialize $A_{k,\delta}(\widetilde{x})\leftarrow\emptyset$.
11:     Initialize cumulative TVD, $D_{TV}\leftarrow 0$.
12:     for each token $x\in B_{k}(\widetilde{x})$ (in increasing distance order) do
13:         Compute potential TVD increment $\Delta D_{TV}$ for adding $x$ to $A_{k,\delta}(\widetilde{x})$.
14:         if $D_{TV}+\Delta D_{TV}<\delta$ then
15:              Add $x$ to $A_{k,\delta}(\widetilde{x})$.
16:              Update $D_{TV}\leftarrow D_{TV}+\Delta D_{TV}$.
17:         else
18:              Break.
19:         end if
20:     end for
21:end for
22:Return proximity sets $B_{k}$ (precomputed) and $A_{k,\delta}$ (dynamically calculated).
Appendix E Experimental Results on Other Visual AR Models
Table 6: Actual speed-up, MAL (Mean Accepted Length), FID, CLIP score, Precision / Recall, and HPS v2 for each method on LlamaGen-XL Stage II and Anole.
Method Acceleration ($\uparrow$) Image Quality Metrics
Speed-up MAL FID ($\downarrow$) CLIP score ($\uparrow$) Prec ($\uparrow$) Rec ($\uparrow$) HPS v2 ($\uparrow$)
LlamaGen-XL Stage II 1.00$\times$ 1.00 47.60 0.2939 0.4138 0.5648 23.84
EAGLE-2 (Li et al., 2024a) 0.96$\times$ 1.22 - - - - -
LANTERN ($\delta=0.4$, $k=1000$) 1.64$\times$ 2.24 46.10 0.2925 0.4704 0.5222 23.06
Anole 1.00$\times$ 1.00 20.27 0.3215 0.6552 0.6398 23.52
EAGLE-2 0.73$\times$ 1.10 - - - - -
LANTERN ($\delta=0.5$, $k=100$) 1.17$\times$ 1.83 23.40 0.3186 0.6026 0.6178 22.92
Table 6 presents additional results comparing LANTERN to EAGLE-2 and standard decoding on LlamaGen-XL Stage II (Sun et al., 2024) and Anole (Chern et al., 2024). The table reports both actual speed-up (measured on RTX 3090 for LlamaGen and A100 80GB SXM for Anole) and image quality metrics, including FID, CLIP score, Precision/Recall, and HPS v2.
As shown in the Table 6, LANTERN consistently achieves superior acceleration compared to EAGLE-2 across multiple visual AR models, including LlamaGen-XL Stage II and Anole. While the performance of LANTERN varies slightly depending on the model, it consistently outperforms EAGLE-2 in terms of both speed-up and mean accepted length. For instance, on the LlamaGen-XL Stage II model, LANTERN achieves a speed-up of 1.64×1.64$\times$1.64 × and a mean accepted length of $2.24$, significantly higher than EAGLE-2’s 0.96×0.96$\times$0.96 × speed-up and $1.22$ mean accepted length. Similarly, for Anole, LANTERN achieves a speed-up of 1.17×1.17$\times$1.17 × with a mean accepted length of $1.83$, compared to EAGLE-2’s 0.73×0.73$\times$0.73 × speed-up and $1.10$ mean accepted length.
In terms of image quality, LANTERN demonstrates minimal degradation despite the improved acceleration. For the LlamaGen-XL Stage II model, the FID decreases slightly to $46.10$ compared to the baseline of $47.60$, while the CLIP score remains stable. Precision and recall metrics indicate that LANTERN maintains a balance between individual image quality (precision) and diversity (recall). On Anole, LANTERN similarly exhibits competitive image quality metrics, with a CLIP score of $0.3186$ and a moderate decrease in HPS v2.
These results highlight the versatility of LANTERN, showcasing its ability to provide significant acceleration benefits while maintaining acceptable image quality across diverse models. This reinforces LANTERN’s effectiveness as a practical and efficient approach for visual AR models.
Appendix F Detailed Latency Analysis for LANTERN
F.1 Analysis on the Number of Captions
Table 7: Actual speedup results for LANTERN across different settings of $\tau$, $k$, and $\delta$, with varying numbers of captions.
Num Actual Speed-up Actual Speed-up Actual Speed-up Actual Speed-up
Captions ($\tau=0$, $k=1000$, $\delta=0.05$) ($\tau=0$, $k=1000$, $\delta=0.2$) ($\tau=1$, $k=1000$, $\delta=0.1$) ($\tau=1$, $k=1000$, $\delta=0.4$)
100 1.56$\times$ 2.33$\times$ 1.13$\times$ 1.73$\times$
1000 1.56$\times$ 2.26$\times$ 1.13$\times$ 1.69$\times$
2000 1.57$\times$ 2.27$\times$ 1.13$\times$ 1.69$\times$
5000 1.56$\times$ 2.26$\times$ 1.13$\times$ 1.69$\times$
Table 7 provides a statistical analysis of actual speed-up measurements conducted with different numbers of captions for MS-COCO validation captions (Lin et al., 2014) on LlamaGen-XL Stage I model (Sun et al., 2024). Across all configurations of $\tau,k$ and $\delta$, the actual speed-ups remain consistent beyond 1000 captions, with negligible differences observed. For example, for $\tau=0,k=1000,\delta=0.05$, the speed-up values remain between 1.56×1.56$\times$1.56 × and 1.57×1.57$\times$1.57 × across 1000, 2000, and 5000 captions. Similarly, other configurations also exhibit minimal variation, confirming the reliability of using 1000 captions as a sample size.
This statistical stability justifies our choice to use 1000 captions in the main experiments, as it provides an accurate and computationally efficient estimate of actual speed-up without sacrificing reliability.
F.2 Component-Level Latency Analysis for LANTERN
Table 8: Component-level latency analysis for LANTERN. Each component’s latency is averaged over a single draft-and-verify process, with measurements conducted on a single RTX 3090 GPU.
Method Target Forward Drafter Forward Proximity Set $A$ Calculation
LANTERN($\delta=0.1,k=1000$) $3.80\times 10^{-2}$s $1.08\times 10^{-2}$s $1.57\times 10^{-3}$s
LANTERN($\delta=0.4,k=1000$) $1.19\times 10^{-3}$s
In this section, we conducted experiments to evaluate the contribution of each component of LANTERN to overall latency. Specifically, we identified three primary components that significantly impact latency: (1) the target model forward pass, (2) the drafter model forward pass, and (3) the computation of the proximity set $A$. Using the LlamaGen-XL Stage I model, we measured the average latency of each component during a single draft-and-verify process across 1,000 prompts from MS-COCO validation captions (Lin et al., 2014). Experiments is conducted under two settings, $k=1000,\delta=0.1$ and $k=1000,\delta=0.4$, with the results summarized in Table 8.
Among the three key contributors to latency, the proximity set computation, which is introduced by LANTERN, required more time at lower $\delta$ values. This can be attributed to the increased number of rejections at lower $\delta$, necessitating multiple proximity set computations within a single step. However, even in the $k=1000,\delta=0.4$ setting, the latency of proximity set computation is 24×24$\times$24 × smaller than that of the target model forward passes and about 7×7$\times$7 × smaller than the drafter model forward pass.
Overall, while LANTERN introduces additional computational overhead compared to traditional speculative decoding methods, the trade-off results in significantly greater speed-up. Thus, our approach demonstrates superior actual speed-up, establishing its advantage over existing speculative decoding methods.
F.3 Latency Comparison on Probability Distance Metrics
Table 9: Computation time comparison between TVD and JSD in LANTERN. Each component’s latency is averaged over a single draft-and-verify process, with measurements conducted on a single RTX 3090 GPU. The hyperparameters for LANTERN are fixed to $k=1000$, and $\delta=0.4$ for TVD and $\delta=0.2$ for JSD.
Distance Metric Computation Time for Distance Total Computation Time
TVD $1.19\times 10^{-3}$s $4.89\times 10^{-2}$s
JSD $4.03\times 10^{-3}$s $4.92\times 10^{-2}$s
Table 9 provides a comparison of computation times between TVD and JSD, supporting our decision to use TVD as the preferred distance metric. We use LlamaGen-XL Stage I with randomly sampled 1000 captions in MS-COCO validation captions (Lin et al., 2014) for this experiment. While both metrics produce nearly identical results in terms of mean accepted length and image quality, TVD is significantly more computationally efficient. Specifically, the computation time for TVD is $1.19\times 10^{-3}$ seconds, which is approximately three times faster than JSD’s computation time of $4.03\times 10^{-3}$ seconds.
Although the total computation time for each decoding step (including other processes) shows a smaller difference, $4.89\times 10^{-2}$ seconds for TVD and $4.92\times 10^{-2}$ seconds for JSD, this difference accumulates over multiple decoding steps. For large-scale applications or models generating long sequences, this efficiency advantage becomes increasingly significant. Therefore, given the negligible impact on quality metrics and the consistent computational advantage, we choose TVD over JSD as the more practical and efficient metric for our method.
Note that the hyperparameters do not affect the computation time for the distance metric itself but do impact the total computation time. Therefore, we set the values of $\delta$ for TVD and JSD to achieve the same level of mean accepted length.
F.4 Latency Results on Intel Gaudi 2 Accelerator
Table 10: Actual speed-up, MAL (Mean Accepted Length) for each method. $\tau=0$ refers to the greedy decoding and $\tau=1$ refers to the sampling with temperature $1$ for generation. The actual speed-up are measured on a single Intel Gaudi 2 (96GB) accelerator and NVIDIA RTX 3090.
Method $\tau=0$
Acceleration ($\uparrow$)
Speed-up (Intel Gaudi 2) Speed-up (NVIDIA RTX 3090) MAL
LlamaGen-XL Stage I 1.00$\times$ 1.00$\times$ 1.00
EAGLE-2 0.80$\times$ 1.29$\times$ 1.60
LANTERN ($\delta=0.05$, $k=1000$) 1.01$\times$ 1.56$\times$ 2.02
LANTERN ($\delta=0.2$, $k=1000$) 1.32$\times$ 2.26$\times$ 2.89
Method $\tau=1$
Acceleration ($\uparrow$)
Speed-up (Intel Gaudi 2) Speed-up (NVIDIA RTX 3090) MAL
LlamaGen-XL Stage I 1.00$\times$ 1.00$\times$ 1.00
EAGLE-2 0.55$\times$ 0.93$\times$ 1.20
LANTERN ($\delta=0.05$, $k=1000$) 0.74$\times$ 1.13$\times$ 1.75
LANTERN ($\delta=0.2$, $k=1000$) 1.02$\times$ 1.69$\times$ 2.40
To demonstrate the adaptability of LANTERN across different hardware platforms, we extend our evaluation beyond NVIDIA GPUs to include the Intel Gaudi 2 accelerator (96GB). Table 10 presents the actual speed-up and Mean Accepted Length (MAL) for different methods, comparing performance on both Intel Gaudi 2 and NVIDIA RTX 3090. While the primary results in the main paper focus on CUDA-based hardware, this analysis highlights that LANTERN can effectively accelerate auto-regressive generation on Gaudi 2 as well.
The observed difference in speed-up between Gaudi 2 and NVIDIA GPUs is not indicative of a general performance gap but rather reflects the unique computational characteristics of speculative decoding. Unlike standard inference and training workloads, where Gaudi 2 has demonstrated competitive performance, speculative decoding introduces irregular execution patterns, including interleaving model forwards and non-trivial memory access patterns. These factors make speculative decoding particularly sensitive to low-level implementation details, which can vary across different accelerator architectures.
Despite the absence of Gaudi-specific optimizations in our implementation, LANTERN still achieves meaningful acceleration on this hardware. While EAGLE-2 achieves only 0.80$\times$ speed-up at $\tau=0$ and 0.55$\times$ at $\tau=1$, LANTERN improves these figures to 1.32$\times$ and 1.02$\times$, respectively. This demonstrates that speculative decoding can be effectively deployed on Gaudi 2 and that the core principles of LANTERN remain beneficial across diverse hardware platforms.
These results suggest that Gaudi 2 is well-suited for auto-regressive generation, including techniques such as speculative decoding. While hardware-specific optimizations could further enhance performance, the existing results already indicate that speculative decoding can be efficiently executed on Gaudi 2. As speculative decoding continues to gain traction in large-scale model deployment, further refinement of its implementation across different accelerator architectures will help maximize its benefits in diverse hardware environments.
Appendix G Size of Latent Proximity Sets Across Positions
(a) $\delta=0.1,k=1000$
(b) $\delta=0.4,k=1000$
Figure 7: Average size of $A_{k,\delta}$ on different position. Note that first token is generated right after pre-fill stage, we do not conduct speculation for first token.
We conduct experiments to better understand the behavior of LANTERN by examining how the size of the latent proximity set $A$ varies depending on the position in images. For $k=1000$ and $\delta=0.1,0.4$, we generate 100 images based on MS-COCO validation captions (Lin et al., 2014) with the LlamaGen-XL Stage I model for each setting. The calculated average size of $A$ on the different positions can be found in Figure 7.
As expected, larger $\delta$ values generally led to a larger size of $A$ regardless of position. Across various $\delta$ values, the size of the proximity set $A$ is consistently larger at the left edge of the image. We hypothesize that this phenomenon occurs because the left edge of the image corresponds to positions immediately following a line break, where uncertainty is higher compared to other positions. This increased uncertainty results in generally lower probabilities assigned to individual tokens, allowing a larger number of tokens to meet the threshold and be included in the set $A$.
Appendix H Trade-Offs with Additional Metrics
Table 11: Precision and recall (Kynkäänniemi et al., 2019) values for varying $\delta$ and $k$ settings. The vanilla AR decoding score is also provided for reference.
Precision / Recall $\tau=0$
$\bm{\delta=0.05}$ $\bm{\delta=0.1}$ $\bm{\delta=0.2}$ $\bm{\delta=0.4}$
Vanilla AR 0.4232 / 0.3517
$\bm{k=100}$ 0.4424 / 0.3212 0.4510 / 0.2989 0.4661 / 0.2754 0.4689 / 0.2813
$\bm{k=300}$ 0.4488 / 0.3195 0.4619 / 0.2960 0.4696 / 0.2802 0.4750 / 0.2753
$\bm{k=1000}$ 0.4484 / 0.3158 0.4659 / 0.2939 0.4771 / 0.2773 0.4854 / 0.2682
Precision / Recall $\tau=1$
$\bm{\delta=0.05}$ $\bm{\delta=0.1}$ $\bm{\delta=0.2}$ $\bm{\delta=0.4}$
Vanilla AR 0.4781 / 0.5633
$\bm{k=100}$ 0.4867 / 0.5389 0.4796 / 0.5303 0.4789 / 0.5140 0.4825 / 0.4946
$\bm{k=300}$ 0.4856 / 0.5367 0.4834 / 0.5231 0.4894 / 0.4901 0.4895 / 0.4719
$\bm{k=1000}$ 0.4865 / 0.5334 0.4869 / 0.5172 0.4880 / 0.4888 0.4909 / 0.4497
Table 12: HPS v2 (Wu et al., 2023) score for different $\delta$ and $k$ settings. The baseline score for Vanilla AR decoding is provided for reference.
HPS v2 $\tau=0$
$\bm{\delta=0.05}$ $\bm{\delta=0.1}$ $\bm{\delta=0.2}$ $\bm{\delta=0.4}$
Vanilla AR 23.18
$\bm{k=100}$ 22.73 22.41 22.13 21.96
$\bm{k=300}$ 22.66 22.25 21.92 21.69
$\bm{k=1000}$ 22.62 22.14 21.69 21.39
HPS v2 $\tau=1$
$\bm{\delta=0.05}$ $\bm{\delta=0.1}$ $\bm{\delta=0.2}$ $\bm{\delta=0.4}$
Vanilla AR 24.11
$\bm{k=100}$ 24.01 23.94 23.86 23.75
$\bm{k=300}$ 23.97 23.85 23.70 23.55
$\bm{k=1000}$ 23.91 23.75 23.47 23.22
To gain a deeper understanding of the trade-offs in LANTERN, we examined precision & recall (Kynkäänniemi et al., 2019), and HPS v2 (Wu et al., 2023) in addition to FID (Heusel et al., 2017), and the results are summarized in Table 11 and 12. The experimental settings are identical to the main experiments in Table 2 and Figure 5.
The result in Table 11 shows that LANTERN achieves comparable or slightly improved precision relative to the baseline across various settings, reflecting its ability to maintain high-quality token generation. While recall decreases marginally with increasing $\delta$, this is an expected trade-off due to the relaxed acceptance condition, which prioritizes sampling speed. The precision/recall results demonstrate that our method strikes a reasonable balance between quality and diversity across different hyperparameter configurations.
To further quantify the aesthetic quality of generated images, we evaluate our approach using HPS v2. Table 12 shows that while there is a slight reduction in aesthetic quality compared to the baseline, the trade-off is well-justified by the significant improvements in generation speed. This aligns with the intended design of LANTERN, which emphasizes efficiency while preserving acceptable quality.
Appendix I Text Prompts for Qualitative Samples
The prompts listed below were used to generate Figure 4, with the images in the figure generated sequentially from left to right using Prompt 1 through Prompt 9.
Prompt 1 A serene lake reflecting the colors of a sunset sky with distant mountains and a few birds flying across the sky.
Prompt 2 A cozy bedroom with soft, warm lighting, a bed with fluffy pillows, a small bookshelf filled with books, and a potted plant on the windowsill.
Prompt 3 A beautiful stained glass window with sunlight streaming through in a church, casting colorful patterns on the stone floor and pews.
Prompt 4 A close-up of a single pink rose in full bloom, with soft, layered petals and gentle sunlight illuminating its delicate curves.
Prompt 5 A classic still life of a bowl of fresh fruit, including apples, oranges, and grapes, with light softly highlighting the textures of each fruit.
Prompt 6 A high-fashion portrait of a woman with vibrant, artistic makeup and bold accessories, wearing a modern, avant-garde outfit against a plain studio background.
Prompt 7 A close-up of a man gazing thoughtfully out of a rain-covered window, with soft reflections and water droplets creating a melancholic, introspective atmosphere.
Prompt 8 A close-up of a small bird perched on a snow-covered branch, with soft, fluffy feathers and a delicate beak, illuminated by gentle winter sunlight.
Prompt 9 A close-up of a wolf with intense, focused eyes and thick gray fur, staring directly at the camera, set against a blurred forest background.
Appendix J Qualitative Results
@@ Appendix I Text Prompts for Qualitative Samples
Figure 9: Qualitative sample for the changes in the generated images according to various $\delta\in\{0.05,0.1,0.2,0.4\}$ and $k\in\{100,300,1000\}$ at $\tau=0$. The input prompt is ’A motorcycle parked in a parking space next to another motorcycle.’. The target model is LlamaGen-XL Stage I.
Figure 10: Additional qualitative samples with various $k$ and $\delta$ at $\tau=1$. The target model is LlamaGen-XL Stage I, and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.
Figure 11: Additional qualitative samples with various $k$ and $\delta$ at $\tau=0$. The target model is LlamaGen-XL Stage I, and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.
Figure 12: Qualitative sample for the changes in the generated images according to various $\delta\in\{0.05,0.1,0.2,0.4\}$ and $k\in\{100,300,1000\}$ at $\tau=1$. The input prompt is ’A pile of oranges in crates topped with yellow bananas.’ The target model is LlamaGen-XL Stage II.
Figure 13: Additional qualitative samples with various $k$ and $\delta$ at $\tau=1$. The target model is LlamaGen-XL Stage II, and MS-COCO validation captions are used. Images within the same column are generated using the same text prompt.